OpenStudy (anonymous):

Sarah went to an electronic store, purchased 3 CDs and 5 DVDs, and spent $137. Michelle went to the same store, bought 2 CDs and 7 DVDs, and spent $161. How much does each CD sell for at the store?

OpenStudy (anonymous):

$13 $14 $18 $19

OpenStudy (anonymous):

B is the answer

OpenStudy (anonymous):

3x + 5y = 137 2x + 7y = 161 6x + 10y = 274 -6x - 21y = -483 -11y = -209 y = 19 x = (274-190)/6 = 14 dollars